consider a spherical shell of radius R at temparature T.  The black body radiation inside it can be considered as an ideal gas of photon with internal energy per unit volume u=u/V is directly proportional to Tand pressure P=1/3×U/V .  If the shell now undergoes an adiabatic expansion the relation between T and  R is 

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
P = U/3V
nRT/V = U/3V
 
nRT/V = T4 /3
 
4/3 pi R^3 T^3 = constant
 
hence T \alpha 1/R
 
Hope it helps
 
Regards
Arun